### v1

# Analysis of Hierarchical AoII over unreliable channel: A Stochastic Hybrid System Approach

Abstract
In this work, we generalize the Stochastic Hybrid Systems (SHSs) analysis of traditional AoI to the AoII metric. Hierarchical ageing processes are adopted using the continuous AoII for the first time, where two different hierarchy schemes, i.e., a hybrid of linear ageing processes with different slopes and a hybrid of linear and quadratic ageing processes, are considered. We first modify the main result in [1, Theorem 1] to provide a systematic way to analyze the continuous hierarchical AoII over unslotted real-time systems. The closed-form expressions of average hierarchical AoII are obtained based on our Theorem 1 in two typical scenarios with different channel conditions, i.e., an M/M/1/1 queue over noisy channel and two M/M/1/1 queues over collision channel. Moreover, we analyze the stability conditions for two scenarios given that the quadratic ageing process may lead to the absence of stationary solutions. Finally, we compare the average age performance between the classic AoI results and our AoII results in the M/M/1/1 queue, and the effects of different channel parameters on AoII are also evaluated.
Index Terms:
Stochastic hybrid system, continuous AoII, hierarchy scheme, stability analysis.
I Introduction
In real-time status update systems, the study of information freshness over the network layer has been the subject of ongoing interest in the past decade, where a classic measure known as Age of Information (AoI) was proposed in [2, 3, 4]. Since then, extensively researches have focused on two dominant directions in this field: i). the analysis of closed-form average AoI (and its moments) over queueing and other networking systems, e.g., [5, 6, 7, 8, 9, 1, 10, 11]; ii). the optimization of AoI performance that searches for the optimal scheduling policies over different networking systems, e.g., [12, 13, 14, 15, 16, 17, 18]. These seminal works also facilitate the cross-fertilization of AoI and other related fields (e.g., topics in Information Theory), and a comprehensive survey of these works can be found in [19].
However, the major drawback of AoI is that this metric only considers the time aspect of update but is content-agnostic, which may pose a challenge to identify the significance of update currently used in the system. According to the basic definition of AoI, the ageing of an update starts since it is generated and continues until it is replaced by the next successful delivered update. However, this kind of ageing measurement may not suitable for all the real-time update system. For example, if the newly generated update contains the same information as the one stored at the monitor, the ageing process of the system should be “paused” for a while since the current update is still fresh enough to be used. Therefore, a variety of modified age metrics have been proposed to address this essential problem, most of which jointly consider the content-mismatch between the transmitter and monitor by introducing suitable error functions as well as the ageing process measured by AoI and its functionals. To our best knowledge, these new metrics include Effective AoI (EAoI) [20], Age of Synchronization (AoS)[21], Query AoI (QAoI)[22], binary freshness[23], Urgency of Information (UoI)[24], and most importantly, Age of Incorrect Information [^1] There are also several variations of AoII, e.g., Age of Changed Information (AoCI)[25], Age of Incorrect Estimates(AoIE)[26], Age of Outdated Information (Ao2I)[27], Age of Processed Information (AoPI)[28]. (AoII) [29]. In fact, it is not difficult to conclude that these metrics could be classified into a more generalized form:
$x(t)=F\Big{(}f(\Delta(t)),g(X(t),\hat{X}(t))\Big{)},$ (1)
where $f:[0,\infty)\to[0,\infty)$ represents the level of dissatisfaction for freshness and $g:\Omega\times\Omega\to[0,\infty)$ defines the error metric function that penalizes the content-mismatch between the stored update and the fresh one at source. Moreover, the mapping $F:[0,\infty)\times[0,\infty)\to[0,\infty)$ describes the relationship between two metrics, e.g., the product form taken in [29]. Certainly, one can select more complex forms such as the exponents when one of the metrics plays a decisive role in achieving a particular communication goal. Hence, following [29], we call the metric in (1) generalized AoII that chooses these mappings to capture the timely utility with three degrees of freedom, which enriches the traditional AoI metric.
Recently, there have been a growing interest in studying the age performance and optimal scheduling policy using AoII, e.g., [30, 31, 32, 33, 34, 38, 35, 36, 37]. As a step in this direction, we are interested in deriving and analyzing the closed-form average age performance by exploring a more general form of AoII using (1). Our work in this paper differs from these aforementioned works in two essential ways:
1) We extend the AoII metric into unslotted real-time update systems using a continuous form of (1), which, to our best of knowledge, has only been used in [37]. Furthermore, we seek for the closed-form expressions of AoII instead of an optimal policy by first introducing Stochastic Hybrid Systems (SHSs) into the analysis of content-aware ageing process, which is essentially different from the analysis in [38] of the coded status update system under different ARQ schemes.
2) We first consider hierarchical ageing processes in different system states since the level of dissatisfaction for freshness would not always grow at a constant rate in practice. On the other hand, since the age and error function are deeply coupled in (1) by the mapping $F$, the penalty of content-mismatch measured by the monitor could be severer depending on the system states (see Fig. 2), which is also reflected in our hierarchical AoII metric.
To summarize, our main contributions in this work are:
$\bullet$ We give the closed-form expressions of average AoII in two typical real-time scenarios: 1) an M/M/1/1 queue over noisy channel; 2) two M/M/1/1 queues over collision channel. The system diagrams are depicted in Fig. 1a and 1b, respectively.
$\bullet$ We provide a systematic way to analyze the continuous hierarchical AoII over unslotted real-time systems using SHSs, where two different hierarchies are utilized. The first is a hybrid of linear functions with different slopes, and the second is a hybrid of linear and quadratic functions. The SHSs approach is more powerful than the traditional renewal-reward approach since the average AoII with quadratic forms cannot be easily obtained by calculating the time averages.
$\bullet$ Considering the potential instability in quadratic ageing, we give our stability conditions in Theorem 2 and 3 for two different systems to ensure the existence of stationary average AoII in terms of the second hierarchy scheme.
$\bullet$ The average age performance over the M/M/1/1 queue measured by AoII is compared with several classic results in traditional AoI in Section V-A. A few insights are obtained based on the inherent differences between AoI and AoII: i). the ceasing of age growth using AoII while the content is matched sharply cuts down the average age in the M/M/1/1 queue compared to AoI results; ii). the average age using AoII can even be smaller than a generate-at-will queue with the same Poisson service rate for some content-change probability $p$; iii). the monotonicity of the average age with the normalized utilization factor $\rho=\frac{\lambda}{\mu}$ is different between two metrics.
$\bullet$ The average AoII performance under different channel parameters in two scenarios are evaluated. In the noisy channel, the average AoII grows monotonically with $p_{e}$ regardless of utilization factors $\rho$ and content-changed probabilities $p$ and the range of $p_{e}$ is limited due to the stability issue. Meanwhile, the average AoII grows monotonically with larger $\rho$ in the collision channel, since a higher arrival rate of any source shall bring a higher collision probability where the mismatch cannot be eliminated and the age keeps growing.
$\bullet$ The computation complexity using Theorem 1 to obtain more general results of multi-access networks is discussed in Section VI, where the number of discrete states grows at least four times the node number $M$.
II System Model
We first give basic assumptions for the noisy channel case with single source and the collision channel case with two sources. Then we present a brief introduction of SHSs and relevant settings for our models.
II-A Model Assumptions
II-A1 The Splitting Poisson Process
We track the average age performance of a source node observing a specific information process $X(t)$. The node can be viewed as a sensor in IoT networks or an observing unit in vehicular networks. The updates $\{X_{t}:t\in\mathbb{R}\}$ are sampled from $X(t)$ following a Poisson process with rate $\lambda$. In this work, we consider a memoryless source model, where each newly sampled update $X_{t}$ would change its content with probability $p$ and remain unchanged with probability $q=1-p$. Consequently, the sampling (arriving) process at the transmitter splits into two Poisson streams with rates $p\lambda$ and $q\lambda$.
The contender source in the collision case is denoted as $X^{c}(t)$, whose updates $\{X^{c}_{t}:t\in\mathbb{R}\}$ are sampled by a node at a Poisson rate $\lambda_{c}$. We assume that the content of $X_{t}$ is independent of the content of $X^{c}_{t}$, therefore, the age behavior of $X(t)$ is only related to whether or not a collision arises.
II-A2 Transmission Through Unreliable Channels
Since time is unslotted, a transmission immediately starts when a newly sampled update arrives. To avoid a combinatorial explosion of the state space, we first assume the transmission times of the updates are modeled as independent exponential $\mu$ and $\mu_{c}$ random variables for updates $X_{t}$ and $X^{c}_{t}$, respectively. Furthermore, the transmission of updates is preemptive considering that the content-changed updates should be delivered as soon as possible, and the preempted update would be discarded. Hence, the transmission at each node follows M/M/1/1.
The channel is unreliable in both cases. In Fig. 1a, an update may be decoded incorrectly by the monitor due to channel noise. We denote $p_{c}$ as the probability of an update being correctly decoded, and an update that is not decoded correctly is lost with probability $p_{e}=1-p_{c}$. Note that only a content-changed update $X_{t}$ being correctly decoded can reduce the age at monitor.
In Fig. 1b, the channel is assumed to be noise-free yet suffering from collision. A collision period begins if a newly sampled update $X^{c}_{t}$ / $X_{t}$ arrives while there is an update from $X_{t}$ / $X^{c}_{t}$ currently being transmitted. The collision disrupts transmission in the way that neither of updates can be successfully decoded by the monitor. The update finishing transmission first would be discarded, whereas the remaining update would keep transmitting. The transmission time of the remaining update is still an exponential random variable due to the memoryless property.
II-A3 End-to-End Interaction and Age Measurements
There are two common assumptions shared in real-time status-updating networks: i). the amounts of communication overheads including controlling signals require negligible transmission time; ii). these overheads are transmitted via a dedicated control channel to ensure the signaling free from errors[39]. Based on that, we can further assume that the indicator of content changes at source is immediately known at the monitor, which plays a key role in characterizing the evolution of our hierarchical AoII. To track the age of $X(t)$, we specify the generalized form (1) in the following:
$x(t)=f(t-u_{t})I\{X_{t}\neq\hat{X}_{t}\},\ t\in\mathbb{R},$ (2)
where $\hat{X}_{t}$ is the last successfully decoded update stored at the monitor, and $u_{t}$ denotes the sampling time of $\hat{X}_{t}$. The indicator function $I$ takes values from 0 or 1 if the content of update being transmitted is the same or not as $\hat{X}_{t}$. Furthermore, we assume that $I$ equal to 0 if there is no update on transmission and the content of source is not changed. The age function $f$ can be linear or quadratic depending on the SHSs state.
[…]
Figure 1: Two system models with memoryless sources: (a) an M/M/1/1 queue over noisy channel. (b) two M/M/1/1 queues over collision channel.
II-B A Brief Introduction to SHSs
Following [40], SHSs are defined by a set of stochastic differential equations
$\mathbf{\dot{x}}=\frac{\partial\mathbf{x}}{\partial t}=f^{\prime}(\mathbf{q},\mathbf{x},t)+g^{\prime}(\mathbf{q},\mathbf{x},t)\frac{\partial\mathbf{z}}{\partial t},$ (3)
where the system states are partitioned into the discrete state $\mathbf{q}\in\mathbb{Q}$ consistent with the ordinary jump process, and a $\mathit{k}$-dimensional stochastic process $\mathbf{x}$ with piecewise continuous sample paths called the continuous state. [^2] For notation simplicity, we here omit the parameter $t$ associated with the joint system state $(\mathbf{q}(t),\mathbf{x}(t))$ at time $t$. Given a discrete state space $\mathbb{Q}$ and a $\mathit{k}$-dimensional vector $\mathbf{z}$ of independent Brownian motion processes, the continuous state $\mathbf{x}$ evolves with two state-dependent functions $f^{\prime}$ and $g^{\prime}$. Moreover, there is a family of discrete transition maps that tracks possible changes of the joint system state $(\mathbf{q},\mathbf{x})$, i.e.,
$(\mathbf{q}^{+},\mathbf{x}^{+})=\phi_{\mathit{l}}(\mathbf{q},\mathbf{x},t),\ \phi_{\mathit{l}}:\mathbb{Q}\times\mathbb{R}^{n}\times[0,\infty)\to\mathbb{Q}\times\mathbb{R}^{n},$ (4)
where $(\mathbf{q}^{+},\mathbf{x}^{+})$ denotes the joint state right after the transition map $\phi_{\mathit{l}}$ takes place, and $\mathit{l}\in\mathcal{L}=\{1,2,...,m\}$ denotes a set of transitions with transition intensities $\lambda_{\mathit{l}}(\mathbf{q},\mathbf{x})$.
Based on the general settings of SHSs above, the continuous state can be defined as the age processes for our AoII models, i.e., $\mathbf{x}(t)=[x_{1}(t),x_{2}(t)]$. We denote $x_{2}(t)$ as the AoII of $X_{t}$ at the monitor, and $x_{1}(t)$ is what the AoII at the monitor would become if an update in service were to complete transmission at time $t$. Since the AoII evolves deterministically at each state $\mathbf{q}$, we have $g^{\prime}(\mathbf{q},\mathbf{x},t)=0$ and $\mathbf{z}=0$. Moreover, the transition rates $\lambda_{l}$ of SHSs are all constant Poisson rates defined above according to the M/M/1/1 queue discipline.
Above all, the hierarchy of AoII of $X_{t}$ can be demonstrated by a piecewise age function $f$, whose derivative is denoted as $f^{\prime}$ in (4). The discrete state $\mathbf{q}$ of SHSs is the system state. Both would be later specified for each case.
$t$$x_{2}(t)$$1_{m}$$1_{m}$$1_{u}$$0_{e}$$0_{c}$$t_{0_{c}}$
$1_{m}$$1_{m}$$0_{c}$$1_{u}$$0_{e}$$t$$x_{2}(t)$$t_{0_{c}}$
Figure 2: The evolution of AoII $x_{2}(t)$ at monitor in the noisy channel case. (a) the linear age penalty with changes of slope in states $1_{u}$ and $0_{e}$; (b) the quadratic age penalty in states $1_{u}$ and $0_{e}$.
III SHSs Analysis over Noisy Channel: The M/M/1/1 queue revisited
The AoI of M/M/1/1 queue has been thoroughly studied under different queueing disciplines. e.g., [5, 41, 1]. However, the queue has never been examined using a context-aware hierarchical age process. In this section, we provide a systematic way to obtain the closed-form results of the average hierarchical AoII in (2) using SHSs. The results would be later compared with the average AoI in Section V.
III-A SHSs Formulations with Hierarchical AoII
To begin, we first need to characterize the system state, i.e., $\mathbf{q}\in\mathbb{Q}=\{0_{c},0_{e},1_{m},1_{u}\}$. State $0_{c}$ is defined as the correct (content-matched) state, where the AoII would not increase hence $\mathbf{\dot{x}}=[0,0]=\mathbf{0}_{n}$. State $1_{m}$ is defined as the mismatched state, where a content-changed update arrives and starts transmission. The AoII $\mathbf{x}(t)$ grows as the traditional AoI in this state, i.e., $\mathbf{\dot{x}}=[1,1]=\mathbf{1}_{n}$. State $0_{e}$ is defined as the error state, where no update is on transmission yet the content between ends is mismatched due to the channel noise. Hence, the age processes $\mathbf{x}(t)$ keep growing in $0_{e}$. State $1_{u}$ is the urgent state, where the transmitter is delivering a more important update. According to the model assumptions, $1_{u}$ can be reached when:
i). a newly sampled update with changed content preempts transmission. In that case, the content of $\hat{X}_{t}$ would lag at least two versions behind the transmitting update;
ii). a newly sampled update recovers the system from the error state. In that case, the monitor needs the update to fix the error as soon as possible.
The hierarchy of (2) can be implemented by adopting different age functions $f$ in each state, where the level of age dissatisfaction at states $0_{e}$ and $1_{u}$ is considered to be severer for the following two reasons.
$\bullet$ The content mismatch between ends cannot be directly eliminated from the error state $0_{e}$ since there is no update on transmission. Being away from $0_{c}$ compared to the other states exacerbates the ageing of $\hat{X}_{t}$ at the monitor.
$\bullet$ The content of update being transmitting at $1_{u}$ is urged by the monitor due to conditions i) and ii) above, which also exacerbates the ageing of $\hat{X}_{t}$.
Consequently, we construct hierarchy schemes with four levels through the state-dependent derivative $f^{\prime}$ using two distinct piecewise functions.
- Linear hierarchy:
$f_{1}^{\prime}(\mathbf{q},\mathbf{x},t)=\begin{cases}\mathbf{0}_{n},\quad\textrm{if}\ \mathbf{q}=0_{c},\\ \mathbf{1}_{n},\quad\textrm{if}\ \mathbf{q}=1_{m},\\ m_{1}\mathbf{1}_{n},\quad\textrm{if}\ \mathbf{q}=0_{e},\\ m_{2}\mathbf{1}_{n},\quad\textrm{if}\ \mathbf{q}=1_{u}.\\ \end{cases}$ (5)
The parameters $m_{1}\geqslant 1$ and $m_{2}\geqslant 1$ are slopes representing two different levels of age dissatisfaction in the linear case.
- Quadratic hierarchy:
$f_{2}^{\prime}(\mathbf{q},\mathbf{x},t)=\begin{cases}\mathbf{0}_{n},\quad\textrm{if}\ \mathbf{q}=0_{c},\\ \mathbf{1}_{n},\quad\textrm{if}\ \mathbf{q}=1_{m},\\ k_{1}\mathbf{x},\quad\textrm{if}\ \mathbf{q}=0_{e},\\ k_{2}\mathbf{x},\quad\textrm{if}\ \mathbf{q}=1_{u}.\\ \end{cases}$ (6)
The parameters $k_{1}$ and $k_{2}$, where $k_{1}<k_{2}$, are the growth constants controlling the stability of SHSs, which would later be constrained by the state transition rates.
$\mathbf{q}=1_{u}$$\mathbf{\dot{x}}=k_{2}\mathbf{x}$$p\lambda$$p_{e}\mu$$\lambda$$\lambda$$p_{c}\mu$$\mathbf{q}=0_{e}$$\mathbf{\dot{x}}=k_{1}\mathbf{x}$$q\lambda$$p_{e}\mu$$\mathbf{q}=1_{m}$$\mathbf{\dot{x}}=\mathbf{1_{n}}$$q\lambda$$p\lambda$$p_{c}\mu$$\mathbf{q}=0_{c}$$\mathbf{\dot{x}}=\mathbf{0}_{n}$
Figure 3: Graphical representation of SHSs: the M/M/1/1 queue with preemption over noisy channel. The hierarchical growth of AoII takes the quadratic case for example, which is distinguished by the color of state. The mappings $\phi_{\mathit{l}}$ is omitted for simplicity.
Remark 1.
The subsequent results and stability analysis are more general using (5) or (6), which can be trivially simplified to a three-level case by letting $m_{1}=m_{2}$ or $k_{1}=k_{2}$. In addition, the hybrid of linear function at state $0_{e}$ and quadratic functions at state $1_{u}$ also follows basically the same analysis procedure as the quadratic case. Therefore, we take (6) as a common example.
To move forward, we now explain each transition $l$ with a constant intensity $\lambda_{\mathit{l}}(\mathbf{q},\mathbf{x})=\lambda_{l}$, which is illustrated in Fig. 3 with the pre-defined SHSs states.
$l=0,3,7$: Since the content of a newly sampled update remains unchanged with probability $q$ in the self-transition at state $0_{c}$, we assume that this kind of update would be maintained at the transmitter subject to an exponential random interval with rate $\mu$ (then dropped) instead of being transmitted. In the self-transition at state $1_{m}$, the preemption of an unchanged update also has no effect on the AoII. Moreover, since we consider the four-level hierarchy in the measure of timeliness for the noisy channel case, the state $1_{u}$ would stay under any preemption while the ageing process reaches the fourth level, which is the highest, at $1_{u}$.
$l=1,5$: An update with changed content leads to the second-level dissatisfaction between ends and the AoII starts to grow after the state is transferred into $1_{m}$ by $l=2$. Similarly, the update under transmission would also be preempted (then discarded) by a newly sampled update with changed content, which leads to the fourth-level dissatisfaction after the state is transferred into $1_{u}$ by $l=5$.
$l=2,6$: The content between ends is correctly matched after a successful transmission over the noisy channel. Meanwhile, the AoII is reset to zero by transitions $2,6$, respectively.
$l=4,8$: If the update is not successfully decoded by the monitor due to noise, the system state would be transferred into the error state $0_{e}$ by $l=4,8$. Then, $\hat{X}_{t}$ at the monitor suffers from a severer age dissatisfaction, namely, the third level in our consideration, due to the mismatch of content.
$l=9$: Any newly sampled update $X_{t}$ starts a transmission in $1_{u}$ when the system was in the error state before. Without loss of generality, we assume that the system goes through different out-of-sync situation between $1_{u}$ and $0_{s}$.
These state transitions are summarized in Table I, along with the possible changes of AoII after mappings $\{\phi_{\mathit{l}}:l\in\mathcal{L}\}$.
TABLE I: State Transitions and Changes of AoII over Noisy Channel
$l$ $\mathbf{q}_{l}\to\mathbf{q}_{l}^{+}$ $\lambda_{l}$ $\mathbf{x}\mathbf{A}_{l}$ $\mathbf{A}_{l}$ $\mathbf{v}_{q_{l}}\mathbf{A}_{l}$
00 $0_{c}\to 0_{c}$ $q\lambda$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
$1$ $0_{c}\to 1_{m}$ $p\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{0_{c}}\end{bmatrix}$
$2$ $1_{m}\to 0_{c}$ $p_{c}\mu$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
$3$ $1_{m}\to 1_{m}$ $q\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{1_{m}}\end{bmatrix}$
$4$ $1_{m}\to 0_{e}$ $p_{e}\mu$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{m}}&v^{2}_{1_{m}}\end{bmatrix}$
$5$ $1_{m}\to 1_{u}$ $p\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{1_{m}}\end{bmatrix}$
$6$ $1_{u}\to 0_{c}$ $p_{c}\mu$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
$7$ $1_{u}\to 1_{u}$ $\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{1_{u}}\end{bmatrix}$
$8$ $1_{u}\to 0_{e}$ $p_{e}\mu$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{u}}&v^{2}_{1_{u}}\end{bmatrix}$
$9$ $0_{e}\to 1_{u}$ $\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{0_{e}}\end{bmatrix}$
III-B Generator for the SHSs
To quantify the average AoII, we first need to employ a test function to track the dynamics of AoII at each state $\overline{\mathbf{q}}\in\mathbb{Q}$,
$\psi_{\overline{\mathbf{q}}}(\mathbf{q},\mathbf{x},t)=\mathbf{x}(t)\delta_{\overline{\mathbf{q}},\mathbf{q}},$ (7)
where $\delta_{\overline{\mathbf{q}},\mathbf{q}}$ is the Kronecker delta function that equals to one only if $\overline{\mathbf{q}}=\mathbf{q}$, otherwise it is zero. We also define
$\mathrm{E}[\psi_{\overline{\mathbf{q}}}(\mathbf{q},\mathbf{x},t)]=\mathrm{E}[\mathbf{x}(t)\delta_{\overline{\mathbf{q}},\mathbf{q}}]=\mathbf{v}_{\overline{\mathbf{q}}}(t),$ (8)
as the expected value of AoII in each state $\overline{\mathbf{q}}$ at time $t$, where $\mathbf{v}_{\overline{\mathbf{q}}}(t)=[v^{1}_{\overline{\mathbf{q}}}(t),v^{2}_{\overline{\mathbf{q}}}(t)]$ tracks two different AoII process. Furthermore, we define the stationary solution (if exists) when $t\to\infty$ to be $\mathbf{v}_{\overline{\mathbf{q}}}=[v^{1}_{\overline{\mathbf{q}}},v^{2}_{\overline{\mathbf{q}}}]=\lim\nolimits_{t\to\infty}\mathbf{v}_{\overline{\mathbf{q}}}(t)$,
Given that $\psi$ is specified and $g(\mathbf{q},\mathbf{x},t)=0$, we introduce one more definition of an operator $L$ on the test function $\psi$[42], that is,
$(L\psi)(\mathbf{q},\mathbf{x},t)=$ $f^{\prime}(\mathbf{q},\mathbf{x},t)\frac{\partial\psi(\mathbf{q},\mathbf{x},t)}{\partial\mathbf{x}}$ (9)
$+\sum_{\mathit{l}=1}^{m}\lambda_{\mathit{l}}\Big{[}\psi(\phi_{\mathit{l}}(\mathbf{q},\mathbf{x},t),t)-\psi(\mathbf{q},\mathbf{x},t)\Big{]},$
which is called the extended generator for our SHSs. $f^{\prime}(\mathbf{q},\mathbf{x},t)$ is in the form of (5) or (6), and $\lambda_{l}$ for the noisy channel case is given in Table I. Note that the extended generator $(L\psi)_{0_{c}}=(L\psi)_{0_{c}}(\mathbf{q},\mathbf{x},t)$ at state $0_{c}$ is zero since the AoII at $0_{c}$ never grows at all.333 As noted in [9], $0_{c}$ is called the irrelevant state to AoII. However, this kind of states does affect the stationary distributions.
With these quantities defined, we are able to introduce the main tool of SHSs in Lemma 1, known as Dynkin formula.
Lemma 1.
Supposing there is a bounded measurable test function $\psi:\mathbb{Q}\times\mathbb{R}^{n}\times[0,\infty)\to\mathbb{R}$, we have
$\frac{\mathrm{d}\mathrm{E}[\psi(\mathbf{q},\mathbf{x},t)]}{\mathrm{d}t}=\mathrm{E}[(L\psi)(\mathbf{q},\mathbf{x},t)].$ (10)
Lemma 1 in fact tells that the dynamics of the test function are on average characterized by its extended generator $L\psi$, whose expected value is the expected rate of $\psi$.
We now can leverage (10) to derive the closed-form expressions of average AoII for the M/M/1/1 queue under linear hierarchy $f_{1}^{\prime}(\mathbf{q})$, which is concluded as [1, Theorem 1]. However, in order to calculate the average AoII under quadratic hierarchy $f_{2}^{\prime}(\mathbf{q})$, we need to modify the aforementioned theorem. Hence, we summarize our first result in the following.
Theorem 1.
Define two sets of transitions
$\mathcal{L}^{i}_{\overline{\mathbf{q}}}=\{l\in\mathcal{L}:\overline{\mathbf{q}}\ \textrm{is the incoming state}\},$ (11)
$\mathcal{L}^{o}_{\overline{\mathbf{q}}}=\{l\in\mathcal{L}:\overline{\mathbf{q}}\ \textrm{is the outgoing state}\},$
as the respective sets of incoming and outgoing transitions for each state $\overline{\mathbf{q}}\in\mathbb{Q}$ associated with $l$. Then, we have:
(i). The finite-state SHSs Markov Chain is ergodic with stationary distributions $\{\pi_{\overline{\mathbf{q}}}:\overline{\mathbf{q}}\in\mathbb{Q}\}$ given by the global balance equations:
$\pi_{\overline{\mathbf{q}}}\sum_{l\in\mathcal{L}^{o}_{\overline{\mathbf{q}}}}\lambda_{l}=\sum_{l\in\mathcal{L}^{i}_{\overline{\mathbf{q}}}}\lambda_{l}\pi_{\mathbf{q}_{l}}.$ (12)
(ii). For linear growth, the expected value $\mathbf{v}_{\overline{\mathbf{q}}}$ can be calculated by
$\mathbf{v}_{\overline{\mathbf{q}}}\sum_{l\in\mathcal{L}^{o}_{\overline{\mathbf{q}}}}\lambda_{l}=f^{\prime}(\overline{\mathbf{q}})\pi_{\overline{\mathbf{q}}}+\sum_{l\in\mathcal{L}^{i}_{\overline{\mathbf{q}}}}\lambda_{l}\mathbf{v}_{\mathbf{q}_{l}}\mathbf{A}_{l},$ (13)
where the age grows as a linear function at $\overline{\mathbf{q}}$, and $\mathbf{q}_{l}$ is denoted as the other state in $\mathcal{L}^{i}_{\overline{\mathbf{q}}}$.
(iii). For quadratic growth, the expected value $\mathbf{v}_{\overline{\mathbf{q}}}$ can be calculated by
$\mathbf{v}_{\overline{\mathbf{q}}}\Big{[}\sum_{l\in\mathcal{L}^{o}_{\overline{\mathbf{q}}}}\lambda_{l}-k_{\overline{\mathbf{q}}}\Big{]}=\sum_{l\in\mathcal{L}^{i}_{\overline{\mathbf{q}}}}\lambda_{l}\mathbf{v}_{\mathbf{q}_{l}}\mathbf{A}_{l},$ (14)
where the age grows as a quadratic function at $\overline{\mathbf{q}}$, and $k_{\overline{\mathbf{q}}}$ is the growth constant associated with $\overline{\mathbf{q}}$.
Consequently, the closed-form expressions of average AoII at the monitor can be obtained by the summation of expected values at each state, i.e.,
$x=\sum_{\overline{\mathbf{q}}\in\mathbb{Q}}\mathbf{v}_{\overline{\mathbf{q}}}^{2},$ (15)
for different functions $f^{\prime}(\overline{\mathbf{q}})$.
The matrix $\mathbf{A}_{l}$ in (13) and (14) is the transition matrix associated with the mapping $\phi_{\mathit{l}}$, which is also given in Table I. Since the first two parts of Theorem 1 have been proved in [1], we omit here for simplicity. The proof of (iii) appears in Appendix A. Note that (14) is only valid under some constrained parameters $k_{1}$ and $k_{2}$, which is called the stability conditions for quadratic hierarchy and is discussed later in Section III-D.
Remark 2.
Theorem 1 gives the general procedure to calculate the average AoII for different hierarchy schemes. First, we need to calculate the stationary distribution for each finite-state SHSs Markov chains using (12), the existence of which ensures the ergodicity. Then, the average AoII can be calculated separately at each state $\mathbf{q}$ using different state-dependent growth function $f^{\prime}(\mathbf{q})$, as derived in (13) and (14). Hence, our theorem can be easily generalized to multi-level hierarchy of state-dependent AoII using any hybrid of linear and quadratic functions. Meanwhile, the theorem can also be further extended to derive the stationary moments (also MGF) of AoII as did in [1, Theorem 1] under the same stability conditions.
III-C Average AoII under Linear Hierarchy
Following Theorem 1, the stationary distributions of the SHSs Markov chain in Fig. 3 are calculated as:
$\pi_{0_{c}}=\frac{p_{c}}{a},$ $\pi_{1_{m}}$ $=\frac{pp_{c}}{a(p+\rho^{-1})},$ (16)
$\pi_{0_{e}}=\frac{pp_{e}}{a},$ $\pi_{1_{u}}$ $=\frac{p^{2}\rho+pp_{e}}{a(p+\rho^{-1})},$
where $a=p+p\rho+p_{c}q$, and $\rho=\lambda/\mu$ is known as the utilization factor of the queue. Then we can use (13) and (15) to obtain the following results.
Corollary 1.
For linear hierarchy scheme (5), the average AoII $x_{a}$ of a preemptive M/M/1/1 queue over the noisy channel at the monitor is
$x_{a}=\frac{m_{1}pp_{e}(1+\rho)}{a\lambda p_{c}}+\frac{(p_{e}+\rho)\big{[}pp_{c}+m_{2}(p^{2}\rho+pp_{e})\big{]}}{a\lambda p_{c}(p+\rho^{-1})},$ (17)
where $m_{1}\geqslant 1$ and $m_{2}\geqslant 1$ are slopes representing two different levels of dissatisfaction for freshness at state $0_{e}$ and $1_{u}$, respectively. Furthermore, if we set $m_{1}=m_{2}=1$, then (17) reduces to the simplest expression with only two levels,
$x^{\prime}_{a}=\frac{p(p_{e}\rho^{-1}+2p_{e}+\rho)}{ap_{c}\mu}.$ (18)
The derivation of (17) appears in Appendix B. The ergodicity of the SHSs Markov chain is established by (16). Hence, it is easy to verify that (17) always exists for any finite $m_{1}$ and $m_{2}$ based on the invertibility of matrix associated with (43), namely, $\det(\mathbf{B})=1-p_{e}>0$ for matrix $\mathbf{B}$ defined in (43). [^4] A general proof can be found in [9, Section V]. The simplified version (18) would later be used to compare with the AoI results in Section V.
Moreover, we note that (17) and (18) can be further specified to the results under metric AoS proposed in [21] by setting $p=1$. To be clearer, since the AoS metric is defined as the time since the content between two ends are out-of-sync, it is exactly the case of the proposed AoII metric where every newly sampled update changes its content with probability one. Hence, the expressions of average AoS in a preemptive M/M/1/1 queue over the noisy channel are given by
$x_{a,AoS}$ $=\frac{m_{1}p_{e}}{\lambda p_{c}}+\frac{p_{c}(p_{e}+\rho)+m_{2}(p_{e}+\rho)^{2}}{\lambda p_{c}(1+\rho)(1+\rho^{-1})},$ (19)
$x^{\prime}_{a,AoS}$ $=\frac{p_{e}\rho^{-1}+2p_{e}+\rho}{p_{c}\mu(1+\rho)},\ \textrm{for}\ m_{1}=m_{2}.$
III-D The Quadratic Hierarchy and Stability Conditions
The average AoII under the quadratic hierarchy is more complex since it involves a hybrid of two kinds of differential equations. The closed-form expression is summarized below.
Corollary 2.
For quadratic hierarchy scheme (6), the average AoII $x_{b}$ of a preemptive M/M/1/1 queue over the noisy channel at the monitor is
$x_{b}=\frac{pp_{c}(p_{e}\mu+\lambda-k_{1})(p\lambda+\mu-k_{2})}{a\lambda(p+\rho^{-1})^{2}\big{[}(\lambda-k_{1})(\mu-k_{2})-\lambda\mu p_{e}\big{]}},$ (20)
under certain stability conditions.
The derivation of (20) appears in Appendix C. However, we are more interested in the stability conditions of (20) since the invertibility of matrix $\mathbf{C}$ defined in (46) depends on the values of $k_{1}$ and $k_{2}$. These conditions are given below, the proof of which can be found in Appendix D.
Theorem 2.
The average hierarchical AoII in (20) exists if and only if the following two conditions are all satisfied:
i). The determinant of the associated matrix $\mathbf{C}$ is large than zero, i.e., $\det(\mathbf{C})=1-\frac{p_{e}\lambda\mu}{(k_{1}-\lambda)(k_{2}-\mu)}>0$;
ii). The growth rates $k_{1}$ and $k_{2}$ at the quadratic growth states is limited, i.e., $0<k_{1}<\lambda<\mu$ and $0<k_{1}<k_{2}<\mu$.
Theorem 2 can be extended to other multi-level (more than four) hierarchies with finer state definitions by going through the same proof procedure as long as the hybrid only consists of linear and quadratic functions. For example, one can split the state $0_{e}$ into two states $0^{1}_{e}$ and $0^{2}_{e}$ representing different levels of error, where $0^{1}_{e}$ is defined the same as the original error state directly connected to $1_{m}$ and $0^{2}_{e}$ is defined as a severer error state that can only be reached by $1_{u}$. Meanwhile, we must point out that the above conditions are actually sufficient and necessary, as proved in Appendix D.
IV SHSs Analysis over Collision Channel
In this section, we focus on another kind of unreliable channel that occurs when multiple nodes attempt to use the channel simultaneously. The SHSs analysis of the collision channel case is basically the same as the noisy channel case, except that the system state space $\mathbb{Q}_{c}$ is larger leading to a much more complex expression of the average AoII. Therefore, we only adopt the quadratic hierarchy scheme in this collision scenario considering the computation similarity and complexity. The stability conditions are also rederived.
IV-A SHSs Formulations with Quadratic Hierarchical AoII
We need to redefine the system states based on the collision nature of channel, where the state space $\mathbb{Q}_{c}$ consists of eight different states $\mathbf{q}\in\mathbb{Q}_{c}=\{0_{c},0_{e},1_{c},1_{m},1_{u},1_{e},2_{m},2_{u}\}$. States $0_{c},0_{e},1_{m}$ and $1_{u}$ are defined the same as the noisy channel case. However, we have to added four more states due to the existence of $X^{c}_{t}$, namely, $1_{c},1_{e},2_{m},2_{u}$. State $1_{c}$ is defined as the other correct (content-matched) state while the contender utilizes the channel to transmit $X^{c}_{t}$ alone. Hence, the AoII in $1_{c}$ would also not increase as in $0_{c}$. On the contrary, state $1_{e}$ is defined as the other error state while the contender utilizes the channel yet the content between ends is mismatched, where we assume that the AoII in $1_{e}$ grows the same as in $0_{e}$. States $2_{m}$ and $2_{u}$ are defined as two distinct collision states under different ageing processes when $X_{t}$ and $X^{c}_{t}$ both occupy the channel. To be clear, the former collision state $2_{m}$ can be directly reached by $1_{m}$ and $1_{c}$, which means the content of $\hat{X}_{t}$ at the monitor only lags one version behind the collision update $X_{t}$. However, the collision update $X_{t}$ in $2_{u}$ is considered to be more urgent for the same two reasons i) and ii) given in III-A, so the ageing process of $\hat{X}_{t}$ is severer than $2_{m}$ at the monitor.
Moreover, since we are only interested in the AoII behavior of $X_{t}$, the continuous states of our SHSs are defined the same as the noisy channel case, i.e., $\mathbf{x}(t)=[x_{1}(t),x_{2}(t)]$ due to the independence between $X_{t}$ and $X^{c}_{t}$.
Based on the above state definitions, the quadratic hierarchy scheme considered here also has four levels, which is characterized by the state-dependent derivative $f_{3}^{\prime}(\mathbf{q},\mathbf{x},t)$.
- Quadratic hierarchy:
$f_{3}^{\prime}(\mathbf{q},\mathbf{x},t)=\begin{cases}\mathbf{0}_{n},\quad\textrm{if}\ \mathbf{q}=0_{c},1_{c},\\ \mathbf{1}_{n},\quad\textrm{if}\ \mathbf{q}=1_{m},2_{m},\\ k_{1}\mathbf{x},\quad\textrm{if}\ \mathbf{q}=0_{e},1_{e},1_{u},\\ k_{2}\mathbf{x},\quad\textrm{if}\ \mathbf{q}=2_{u}.\\ \end{cases}$ (21)
The parameters $k_{1}$ and $k_{2}$ are the growth constants controlling the stability of SHSs, where $k_{1}<k_{2}$. Given that the collision channel is noise-free, state $2_{u}$ is assumed to be the only urgent state with the highest age dissatisfaction (the fourth level) since it is the only state that suffers from both collision and a severer out-of-sync situation. Note that this assumption can be easily generalized by adapting the growth rate constant $k_{1}$ and $k_{2}$.
Remark 3.
As we emphasized in the noisy channel case, the analysis and conclusion in this section could be extended to other multi-level hierarchies with finer state definitions. For example, one can track different urgent states in which successive arrivals of content-changed updates can lead to different levels of age dissatisfaction. However, these extensions are not trivial even in the two node case and each hierarchy scheme needs to be carefully designed due to the computation complexity of the SHSs approach, which can be indicated by our case below.
$\mathbf{q}=1_{u}$$\mathbf{\dot{x}}=k_{1}\mathbf{x}$$\lambda$$\lambda_{c}$$\mu$$\mathbf{q}=1_{m}$$\mathbf{\dot{x}}=\mathbf{1_{n}}$$q\lambda$$\mu$$\lambda_{c}$$p\lambda$$\mathbf{q}=1_{e}$$\mathbf{\dot{x}}=k_{1}\mathbf{x}$$\lambda_{c}$$\lambda$$\mu_{c}$$\mathbf{q}=1_{c}$$\mathbf{\dot{x}}=\mathbf{0_{n}}$$\lambda_{c}+q\lambda$$\mu_{c}$$p\lambda$$\mathbf{q}=2_{u}$$\mathbf{\dot{x}}=k_{2}\mathbf{x}$$\lambda_{c}+\lambda$$\mu_{c}$$\mu$$\mathbf{q}=2_{m}$$\mathbf{\dot{x}}=\mathbf{1_{n}}$$\lambda_{c}+q\lambda$$\mu_{c}$$p\lambda$$\mu$$\mathbf{q}=0_{e}$$\mathbf{\dot{x}}=k_{1}\mathbf{x}$$\lambda$$\lambda_{c}$$\mathbf{q}=0_{c}$$\mathbf{\dot{x}}=\mathbf{0_{n}}$$q\lambda$$p\lambda$$\lambda_{c}$
Figure 4: Graphical representation of SHSs: two M/M/1/1 queues over collision channel. The mappings $\phi_{\mathit{l}}$ is omitted for simplicity.
The SHSs states and corresponding transition set $\mathcal{L}_{c}$ are depicted in Fig. 4. For clarity, we now explain each transition $l\in\mathcal{L}_{c}$.
$l=0,6,9,13,16,20,24$: First, the assumption of the self-transition at state $0_{c}$ remains the same as the noisy channel case. In the remaining self-transitions, states $1_{c}$, $1_{m}$ and $2_{m}$ would stay unchanged when the transmission of updates is preempted either by a content-unchanged update (maintained) $X_{t}$ or the contender $X^{c}_{t}$. Moreover, states $1_{u}$, $1_{e}$ and $2_{u}$ would stay unchanged under any preemption.
TABLE II: State Transitions
$l$ $\mathbf{q}_{l}\to\mathbf{q}_{l}^{+}$ $\lambda^{(l)}$ $\mathbf{x}\mathbf{A}_{l}$ $\mathbf{A}_{l}$ $\mathbf{v}_{q_{l}}\mathbf{A}_{l}$
00 $0_{c}\to 0_{c}$ $q\lambda$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
$1$ $0_{c}\to 1_{c}$ $\lambda_{c}$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
$2$ $0_{c}\to 1_{m}$ $p\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{0_{c}}\end{bmatrix}$
$3$ $0_{e}\to 1_{e}$ $\lambda_{c}$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{0_{e}}&v^{2}_{0_{e}}\end{bmatrix}$
$4$ $0_{e}\to 1_{u}$ $\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{0_{e}}\end{bmatrix}$
$5$ $1_{c}\to 0_{c}$ $\mu_{c}$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
$6$ $1_{c}\to 1_{c}$ $\lambda_{c}+q\lambda$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
$7$ $1_{c}\to 2_{m}$ $p\lambda$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{c}}&v^{2}_{1_{c}}\end{bmatrix}$
$8$ $1_{m}\to 0_{c}$ $\mu$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
$9$ $1_{m}\to 1_{m}$ $q\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{1_{m}}\end{bmatrix}$
$10$ $1_{m}\to 2_{m}$ $\lambda_{c}$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{m}}&v^{2}_{1_{m}}\end{bmatrix}$
$11$ $1_{m}\to 1_{u}$ $p\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{1_{m}}\end{bmatrix}$
$12$ $1_{u}\to 0_{c}$ $\mu$ $\begin{bmatrix}0&0\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&0\end{bmatrix}$ $\begin{bmatrix}0&0\end{bmatrix}$
$13$ $1_{u}\to 1_{u}$ $\lambda$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{1_{u}}\end{bmatrix}$
$14$ $1_{u}\to 2_{u}$ $\lambda_{c}$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{u}}&v^{2}_{1_{u}}\end{bmatrix}$
$15$ $1_{e}\to 0_{e}$ $\mu_{c}$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{e}}&v^{2}_{1_{e}}\end{bmatrix}$
$16$ $1_{e}\to 1_{e}$ $\lambda_{c}$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{e}}&v^{2}_{1_{e}}\end{bmatrix}$
$17$ $1_{e}\to 2_{u}$ $\lambda$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{1_{e}}&v^{2}_{1_{e}}\end{bmatrix}$
$18$ $2_{m}\to 1_{m}$ $\mu_{c}$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{2_{m}}\end{bmatrix}$
$19$ $2_{m}\to 1_{e}$ $\mu$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{2_{m}}&v^{2}_{2_{m}}\end{bmatrix}$
$20$ $2_{m}\to 2_{m}$ $\lambda_{c}+q\lambda$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{2_{m}}&v^{2}_{2_{m}}\end{bmatrix}$
$21$ $2_{m}\to 2_{u}$ $p\lambda$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{2_{m}}&v^{2}_{2_{m}}\end{bmatrix}$
$22$ $2_{u}\to 1_{u}$ $\mu_{c}$ $\begin{bmatrix}0&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$ $\begin{bmatrix}0&v^{2}_{2_{u}}\end{bmatrix}$
$23$ $2_{u}\to 1_{e}$ $\mu$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{2_{u}}&v^{2}_{2_{u}}\end{bmatrix}$
$24$ $2_{u}\to 2_{u}$ $\lambda_{c}+\lambda$ $\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ $\begin{bmatrix}0&0\\ 1&1\end{bmatrix}$ $\begin{bmatrix}v^{2}_{2_{u}}&v^{2}_{2_{u}}\end{bmatrix}$
$\pi^{\prime}_{0_{c}}=\frac{m\mu\mu_{c}(p\lambda+\mu_{c})}{b(\lambda_{c}+\mu_{c})},\ \pi^{\prime}_{1_{c}}=\frac{m\lambda_{c}\mu\mu_{c}}{b(\lambda_{c}+\mu_{c})},\pi^{\prime}_{0_{e}}=\frac{p\lambda_{c}\mu\mu_{c}(m-q\lambda)}{b(\lambda_{c}+\mu_{c})},\ \pi^{\prime}_{1_{e}}=\frac{p\lambda_{c}\mu(\lambda_{c}+\lambda)(m-q\lambda)}{b(\lambda_{c}+\mu_{c})}$ (22a)
$\pi^{\prime}_{1_{m}}=\frac{mp\lambda\mu\mu_{c}\big{[}(p\lambda+\mu+\mu_{c})(p\lambda+\mu_{c})+\lambda_{c}\mu_{c}\big{]}}{b(\lambda_{c}+\mu_{c})(p\lambda+\mu)(m-q\lambda)},\ \pi^{\prime}_{2_{m}}=\frac{mp\lambda_{c}\lambda\mu\mu_{c}(m-q\lambda)+mp^{2}\lambda_{c}\lambda^{2}\mu\mu_{c}}{b(\lambda_{c}+\mu_{c})(p\lambda+\mu)(m-q\lambda)},$
$\pi^{\prime}_{1_{u}}=\frac{mp\lambda\mu_{c}\Big{\{}\big{[}p\lambda(p\lambda+\lambda_{c}+\mu_{c})+\lambda_{c}\mu\big{]}(m-q\lambda)+p\lambda_{c}\lambda\mu\Big{\}}}{b(\lambda_{c}+\mu_{c})(p\lambda+\mu)(m-q\lambda)},\ \pi^{\prime}_{2_{u}}=1-\sum_{\mathbf{q}\in\mathbb{Q}_{c}/\{2_{u}\}}\pi_{\mathbf{q}}.$
$x_{c}=\frac{1}{\Pi}$$\Biggl{\{}c_{2}\lambda^{2}(\lambda_{c}+\mu+\mu_{c}-k_{1})\big{[}p(m-2k_{1})+\mu\big{]}+c_{1}p\lambda^{2}\big{[}(\lambda-k_{1})(\lambda_{c}+\mu_{c}-k_{2})+(\lambda_{c}+\mu_{c}-k_{1})(\lambda_{c}+\mu+\mu_{c}-k_{2})\big{]}$ (22b)
$-c_{2}\lambda\biggl{\{}p(\lambda_{c}+\mu_{c}-k_{1})\big{[}k_{1}(\lambda_{c}+\mu_{c}-k_{1})-\mu(\lambda_{c}-2k_{1})\big{]}-\mu^{3}-\mu(\lambda_{c}+\mu_{c}-k_{1}-k_{2})\big{[}\mu(p+2)+\lambda_{c}+\mu_{c}-k_{1}\big{]}\biggr{\}}$
$-c_{1}p\lambda(\lambda_{c}+\mu_{c}-k_{1})\big{[}\mu(\lambda_{c}-k_{1})-k_{1}(\lambda_{c}+\mu_{c}-k_{2})\big{]}$
$+c_{2}\mu(\lambda_{c}+\mu_{c}-k_{1})\big{[}\mu^{2}+\mu(\lambda_{c}+\mu_{c}-k_{1}-k_{2})-k_{1}(\mu_{c}-k_{2})-k_{2}\lambda_{c}\big{]}\Biggr{\}}+c_{1}+c_{2}.$
$\Pi=\det(\mathbf{D})(\lambda+\mu_{c}-k_{1})(\lambda+\lambda_{c}-k_{1})(\lambda_{c}+\mu-k_{1})(\mu+\mu_{c}-k_{2}).$ (22c)
$l=1,2$: Since a newly sampled update $X^{c}_{t}$ has no effect on the AoII of $X_{t}$ if the contender does not cause collision, the transition $l=1$ keeps $x_{1}=x_{2}=0$. However, A newly sampled update $X_{t}$ with changed content leads to mismatch between ends, and the AoII starts to grow subject to the linear (second level) age dissatisfaction after the state is transferred into $1_{m}$ by $l=2$.
$l=5,8,12$: If there is no collision, the content between ends would be correctly matched and the AoII is set to zero after a successful transmission corresponding to transitions $5,8,12$, respectively.
$l=11,21$: The update $X_{t}$ being transmitted is preempted by a new content-changed update. Hence, the mismatch between ends becomes severer and the states are transferred into $1_{u}$ and $2_{u}$ respectively. Since the preemption has no effect on the channel collision, the transition $l=21$ would keep the continuous state unchanged, i.e., $\mathbf{A}_{l}=\mathbf{I}$ is the identity matrix.
$l=7,10,14,17$: For transition $l=7$, the channel would collide if a newly sampled update $X_{t}$ with changed content arrives when the channel is currently busy at transmitting the update $X^{c}_{t}$ from contender. Similarly, for transitions $l=17$, the channel collides for any newly arrived update $X_{t}$ if $X^{c}_{t}$ is being transmitted. For transitions $l=10,14$, the channel would collide if any newly sampled update of the contender arrives when the channel is currently busy at transmitting the update $X_{t}$. The difference between $l=7,10$ and $l=14,17$ is in that the latter transfer the state into $2_{u}$, which suffers from the fourth level age dissatisfaction due to a severer out-of-sync situation between ends. Since the collision guarantees that neither update in transmission is successfully received, these four transitions also set $\mathbf{x}\mathbf{A}_{l}=\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$.
$l=18,19,22,23$: The channel becomes available from collision if either $X_{t}$ or $X^{c}_{t}$ is unsuccessfully delivered and discarded, where the transmission time of the left update is still exponential. The difference between $l=18,22$ and $l=19,23$ is in that the latter transfer the state into $1_{e}$, where no update of $X_{t}$ is being transmitted yet the mismatch of content still exists. Hence, $l=19,23$ lead to $\mathbf{x}\mathbf{A}_{l}=\begin{bmatrix}x_{2}&x_{2}\end{bmatrix}$ instead of $\mathbf{x}\mathbf{A}_{l}=\begin{bmatrix}0&x_{2}\end{bmatrix}$ in $l=18,22$.
$l=3,15$: The transitions $l=3,15$ between states $0_{e}$ and $1_{e}$ only depend on the contender source, which cannot lead to any possible drops on the AoII of $X_{t}$. Moreover, $x_{2}(t)$ keeps growing at the third level in these two error states because the content mismatch still exists. Hence, the continuous states are unchanged, i.e., $\mathbf{A}_{l}=\mathbf{I}$.
$l=4$: As assumed in the noisy channel case, any newly sampled update $X_{t}$ starts a transmission in $1_{u}$ when the system was in $0_{e}$ before.
The state transitions above are summarized in Table II, along with the possible changes of AoII after mappings $\{\phi_{\mathit{l}}:l\in\mathcal{L}_{c}\}$.
IV-B Average AoII under Quadratic Hierarchy
Following (12), the stationary distributions of the SHSs Markov chain in Fig. 4 are calculated as (LABEL:collision_distribution), where $m=\lambda_{c}+\lambda+\mu_{c}+\mu$ and $b=p(m\lambda+\lambda_{c}\mu)(m-q\lambda)+m\mu\mu_{c}$. [^5] The explicit expression of $\pi_{2_{u}}$ cannot be easily factorized, which is too long to be included in (LABEL:collision_distribution). Hence, we choose to present the implicit form.
Moreover, we could further obtain the average AoII of the collision channel case under a four-level hierarchy using Theorem 1.
Corollary 3.
For quadratic hierarchy scheme (21), the average AoII $x_{c}$ of two M/M/1/1 queues over the collision channel at the monitor is summarized as (22b) under certain stability conditions, where the denominator is given in (22c).
The derivation of Corollary 3 appears in Appendix E, where the constants $c_{1}$ and $c_{2}$ are given in (70) as the expected values of AoII at states $1_{m}$ and $2_{m}$, respectively. The associated matrix $\mathbf{D}$ is defined in (71), where the explicit expression of it determinant $\det(\mathbf{D})$ is given in Appendix F and is omitted here for simplicity.
IV-C Sufficient Conditions of Stability over Collision Channel
The stability conditions of the general case (22b) for any $\lambda,\lambda_{c},\mu,\mu_{c}$ are hard to obtain since the stable growth rates $k_{1}$ and $k_{2}$ depend on the relationship between two pairs $(\lambda,\mu)$ and $(\lambda_{c},\mu_{c})$. Therefore, we now present some sufficient conditions for stability control in collision channel and give our finer results for a specially symmetric case where the contender $X^{c}_{t}$ is evenly matched to $X_{t}$, i.e., $\lambda=\lambda_{c}$ and $\mu=\mu_{c}$.
Theorem 3.
Assuming that the following conditions are satisfied in the collision channel:
i). The determinant of the associated matrix $\mathbf{D}$ is large than zero, i.e., $\det(\mathbf{D})>0$;
ii). The growth rates $k_{1}$ and $k_{2}$ at the quadratic growth states is limited, i.e., $0<k_{1}<\lambda$ and $0<k_{1}<k_{2}<\mu$,
then (22b) exists for any $\lambda,\lambda_{c},\mu,\mu_{c}$.
Furthermore, if $\lambda=\lambda_{c}$ and $\mu=\mu_{c}$, the above conditions can be modified as:
$0<k_{1}<\lambda<\mu,$ (23)
$0<k_{1}<k_{2}<\min\{\mu,\frac{2k_{1}\mu^{2}+4k_{1}\lambda\mu-2k_{1}^{2}\mu-2\lambda\mu^{2}}{(3\lambda+\mu)k_{1}-k_{1}^{2}-\lambda\mu-2\lambda^{2}}\}.$
The proof of Theorem 3 appears in Appendix F. Note that (23) helps to reveal the explicit relationship between $k_{1}$ and $k_{2}$ similar to Theorem 2 with the knowledge of the rates of contender.
Remark 4.
We must admit that the whole range of $k_{1}$ and $k_{2}$ ensuring stability might be (much) larger than the intervals given in Theorem 3, which are affected by some unknown contender. However, the explicit relationship between $k_{1}$ and $k_{2}$ given by condition i) are intractable unless we know the relationship of rates between the contender and the source we interest. Nevertheless, it is sufficient to adopt some appropriate constants indicated by the above conditions to achieve two main goals: a). the ability to distinguish different requirements of freshness at different system states; b). the ability to ensure the stability of real-time stochastic communication systems with limited consideration of the effects from environment. In practice, we prefer to choose the parameters $k_{1}$ and $k_{2}$ smaller than $\lambda$ which also are independent of the rates of contender.
V Numerical Results
In this section, we compare the classical results over the M/M/1/1 queue using traditional AoI and AoII metric. Meanwhile, numerical results of the average hierarchical AoII over noisy channel and collision channel are given under different parameters, respectively. The effects of different channel conditions are also analyzed.
Figure 5: Illustration of the comparisons of the average age over the M/M/1/1 queue between AoI and AoII
Figure 6: The average AoII over noisy channel with changing parameters $p$ and $p_{e}$, where the growth rates $k_{1}=1$ and $k_{2}=1.5$. The curves under different utilization factors $\rho=0.2,0.5,0.9$ are given in (a), (b), (c), respectively.
V-A Comparisons between AoI and AoII: The M/M/1/1 queue
To begin, we first need to introduce several results obtained in the M/M/1/1 queue using AoI, where the parameters $\lambda$ and $\mu$ are arrival and service rates of updates.
1). The M/M/1/1 queue: an update is accepted in the system only if the server is idle, furthermore, new arrivals are discarded when there is an update in service. The average age is given by [5, Equ. (21)], i.e.,
$\Delta_{M/M/1/1}=\frac{1}{\lambda}+\frac{2}{\mu}-\frac{1}{\lambda+\mu}.$ (24)
2). The M/M/1/1 queue with preemption: an update-in-service is preempted and discarded if there is a new arrival. The average age is given by [41, Equ. (8)]666also see [43, Equ. 18]., where the source only generates updates of one kind and the service time is set to be exponential, i.e.,
$\Delta^{p}_{M/M/1/1}=\frac{1}{\lambda}+\frac{1}{\mu}.$ (25)
3). The M/M/1/1 queue with abandonment: an update-in-service is abandoned (i.e. discarded without completing service) at rate $\alpha$. The average age is given by [1, Equ. (24)],
$\Delta^{\alpha}_{M/M/1/1}=\frac{1}{\lambda}+\frac{1}{\mu}+\frac{\lambda}{(\mu+\alpha)(\lambda+\mu+\alpha)}+\frac{\alpha}{\lambda\mu}.$ (26)
Note that if we set the abandonment rate $\alpha=0$, then we have the traditional case in (24).
Since the channel of the above AoI results is assumed to be noise-free, we need to modify our AoII result in (18) for a fair comparison by setting $p_{e}=0$, which results in:
4). The average AoII of the M/M/1/1 queue with preemption:
$x^{\ast}_{a}=\frac{p\lambda}{(p\lambda+\mu)\mu}=\frac{1}{\mu}-\frac{1}{p\lambda+\mu}.$ (27)
Meanwhile, the AoS results over the noisy-free channel can obtained from (27) by letting $p=1$, that is,
5). The average AoS of the M/M/1/1 queue with preemption:
$x^{\ast}_{a,AoS}=\frac{\lambda}{(\lambda+\mu)\mu}=\frac{1}{\mu}-\frac{1}{\lambda+\mu}.$ (28)
We first observe that
$\Delta^{p}_{M/M/1/1}>\frac{2}{\mu},\quad x^{\ast}_{a,AoS}<\frac{1}{2\mu},$ (29)
under the assumption $\lambda<\mu$. Then, we have the relationship in the preemption $M/M/1/1$ queue among three different age metrics, that is,
$\Delta^{p}_{M/M/1/1}>4x^{\ast}_{a,AoS}>4x^{\ast}_{a}.$ (30)
The reason of the first inequality is that the AoI metric incorporates a redundant ageing process when the content of update is unchanged (still fresh) at the transmitter. The reason of the second inequality is that the possible change of content ($p<1$) using AoII prolongs the occupancy time of the correct state $0_{c}$ compared to the AoS case ($p=1$). Since the age at $0_{c}$ is always zero, the prolongation helps to further reduce the average.
Interestingly, we also find that
$\Delta_{M/M/1/1}=\Delta^{p}_{M/M/1/1}+x^{\ast}_{a,AoS}.$ (31)
Combined with (30), we have
$\Delta_{M/M/1/1}>5x^{\ast}_{a,AoS}>5x^{\ast}_{a}.$ (32)
Equations (30) and (32) indicates that the ceasing of age growth at the matched state sharply cuts down the average age in the M/M/1/1 queue systems, which in turn enables a large gain on the system performance of information freshness.
Last but not least, we note that
$x^{\ast}_{a,AoS}<\frac{1}{\mu},$ (33)
where the constant $\frac{1}{\mu}$ at the right-hand side can be interpreted as the average age of a generate-at-will queue with Poisson service rate $\mu$. A generate-at-will queue is the queue that a new content-changed update arrives exactly when the current update finishes transmission. Hence, this inequality reveals that the occupancy time of the state $0_{c}$ in (28) in fact brings a reduction of $\frac{1}{\lambda+\mu}$ in the average age.
Moreover, we can further conclude that the average age relationship between (27) and the generate-at-will queue under the same content-change probability $p$, whose average age is $\frac{p}{\mu}$, depends on $p$ and the utilization factor $\rho$.
We illustrate the above results numerically in Fig. 5 by setting the service rate $\mu=1$, where we adopt the natural logarithm of the average age. Note that a larger abandonment rate $\alpha$ leads to a smaller average AoI, whereas a larger content-change probability $p$ gives a larger average AoII.
More interestingly, the average AoI is monotonically decreasing with the Poisson arrival rate $\lambda$ (equivalently, the utilization factor $\rho$) yet the average AoII is monotonically increasing with $\lambda$. The reason for this paradox lies in the inherent difference between two age metrics. To be specific, a larger arrival rate in the AoII-based M/M/1/1 queue shortens the occupancy time in $0_{c}$, which in turn gives a larger average system age due to the frequently changed content. On the contrary, in the AoI-based M/M/1/1 queue, a larger arrival rate shortens the occupancy time when no update is on transmission yet the AoI still grows, which ensures a smaller average system age.
V-B The effects of noise in the M/M/1/1 queue
The curves of (20) are depicted in Fig. 6 with $k_{1}=1$ and $k_{2}=1.5$. Two main effects of channel noise on the average AoII can be concluded:
1). The average AoII grows monotonically with $p_{e}$ regardless of utilization factors $\rho$ and content-changed probabilities $p$ in Fig. 6. The reason is obvious since the mismatch between ends cannot be timely eliminated due to large $p_{e}$.
2). The range of $p_{e}$ is limited, e.g., $p_{e}<0.4$, $p_{e}<0.6$ and $p_{e}<0.7$ in Fig. 6a, 6b and 6c, respectively, due to the first stability condition given in Theorem 2. Based on this, we could conclude that a poor channel condition may lead to instability even the growth rates are relatively small. In SHSs, it means that any $p_{e}$ violating the stability condition results in a finite escape time of the AoII with non-zero probability.
Furthermore, it is interesting to note that the average AoII at a poor channel condition (for instance, $p_{e}>0.5$) need not always grow monotonically with larger $p$ as seen in Fig. 6c, e.g., the high utilization case. One reasonable guess of this inconsistency could be that the high content-changed probability $p$ leads to relatively small occupancy time in state $0_{e}$ and large occupancy time in state $0_{c}$ when the system state is more likely to transfer into $1_{u}$. Since the difference is not significant between growth rates $k_{1}$ and $k_{2}$, the average age under this circumstance is lower instead.
Figure 7: The average AoII over collision channel in symmetric case, where the growth constants $k_{1}=1$ and $k_{2}=5$ are chosen to satisfy the stability condition in Theorem 3.
V-C The effects of collision in two M/M/1/1 queues
Considering the complexity of (22b), we only illustrate the symmetric case, i.e., $\lambda=\lambda_{c}$ and $\mu=\mu_{c}$, where the stability conditions in Theorem 3 are easy to obtain. The curves of different values of utilization factors (with $\mu=\mu_{c}=10$ fixed) are depicted in Fig. 7 with varying content-changed probability $p$. The average AoII increases monotonically with larger utilization factors as expected since the higher arrival rates of updates from both sources are more likely to cause channel collision.
VI Discussions and Conclusions
In this paper, we consider two different hierarchy schemes, i.e., the linear and quadratic hierarchies, for different discrete states of SHSs, and provide a systematic way to derive the closed-form expressions of average hierarchical AoII. Moreover, we give the stability conditions for the quadratic hierarchy in both channel cases by exploiting the linear systems established by SHSs, which are essential for the existence of the average AoII (also the higher moments).
The derivation of explicit expressions of different age moments in stochastic systems are commonly classified into a set of problems in renewal-reward process. As an alternative of traditional renewal-reward methods, the SHSs method introduced by [9] is considered to be more powerful in that it can not only deal with age-dependent transition rates, which is first studied in [44], also the age-dependent hierarchies that is first considered in our work. However, the limitations of SHSs are also apparent from the collision channel case, i.e., the computation complexity of Theorem 1. To be clearer, one could discuss two simple cases of generalization:
i). variations of multi-level hierarchical AoII with a finer state definitions;
ii). general multi-access cases with $M\geqslant 3$ with a larger finite state space $\mathbb{Q}_{M}$ considering a $k$-way collision.
Both generalizations follow the same analysis procedure, yet the computation cost is enormous due to the order of the associated matrix. Case i) has been discussed after Theorem 2 also in Remark 3. In case ii), the state definition for AoII is more complex than the AoI case in [45]. One should also track the number $k$ of active transmitters as in [45], however, the simplest definition of $\mathbb{Q}_{M}$ is $\mathbb{Q}_{M}=\mathbb{S}_{1}\cup\mathbb{S}_{2}\cup\mathbb{S}_{3}$, where
$\mathbb{S}_{1}$ =$=\{0_{e},0_{c}\},\ \mathbb{S}_{2}=\{M_{m},M_{u}\},$ (34)
$\mathbb{S}_{3}$ =$=\{k_{c},k_{m},k_{u},k_{e}:1\leqslant k\leqslant M-1\}.$
The meanings of states are defined similarly as in the two-node collision case according to their indexes. Hence, the number of states in $\mathbb{Q}_{M}$ is at least $4M$, which in turn gives a set of linear equations with the size of the associated matrix being $3M\times 3M$777the number of irrelevant states is $M$ Although one could solve for the above generalizations with the help of computer software, the expressions could be very complex and no stability conditions could be easily obtained. Based on this, a promising direction on the analysis of continuous AoII in more complex stochastic (queue) systems could be the searches for appropriate bounds.
Stochastic control and policy improvement are not discussed, which are still fresh topics in the AoII literature, especially in the continuous case. Our undergoing works are devoted to address these state-of-the-art problems.
Appendix A Proof of Theorem 1: The Third Part
Since it follows from (7) and Lemma 1 that the change of expected values of $\psi_{\overline{\mathbf{q}}}$ is controlled by an individual stochastic differential equation at each state $\overline{\mathbf{q}}\in\mathbb{Q}$, we only need to rederive the formulas of expected values $\mathbf{v}_{\overline{\mathbf{q}}}$ at quadratic growth states. Without loss of generality, we leverage state $1_{u}$ of the noisy channel case to prove (14).
First, we need to represent the extended generator at ${1_{u}}$ as
$(L\psi)_{1_{u}}(\mathbf{q},\mathbf{x},t)=$ $k_{2}\mathbf{x}\delta_{1_{u},\mathbf{q}}$ (35)
$+\sum_{l=1}^{m}\lambda_{l}\Big{[}\psi_{1_{u}}\big{(}\phi_{\mathit{l}}(\mathbf{q},\mathbf{x},t)\big{)}-\psi_{1_{u}}(\mathbf{q},\mathbf{x},t)\Big{]}.$
The first term in the right-hand side of (35) follows from
$f^{\prime}_{2}(1_{u})\frac{\partial\psi_{1_{u}}(\mathbf{q},\mathbf{x},t)}{\partial\mathbf{x}}$ $=f^{\prime}_{2}(1_{u})\mathbf{I}_{2}\delta_{1_{u},\mathbf{q}}$ (36)
$=k_{2}\mathbf{x}\delta_{1_{u},\mathbf{q}},$
where $\mathbf{I}_{2}$ is the $2\times 2$ identity matrix. The second term in the right-hand side of (35) can be divided into two parts according to (11), that is,
$\sum_{l=1}^{m}\lambda_{l}\Big{[}\psi_{1_{u}}\big{(}\phi_{\mathit{l}}(\mathbf{q},\mathbf{x},t)\big{)}-\psi_{1_{u}}(\mathbf{q},\mathbf{x},t)\Big{]}$ (37)
$=$ $\sum_{l\in\mathcal{L}^{o}_{1_{u}}}\lambda_{l}\delta_{1_{u},\mathbf{q}}\Big{[}\psi_{1_{u}}\big{(}\mathbf{q}^{+},\mathbf{x}^{+}\big{)}-\psi_{1_{u}}(\mathbf{q},\mathbf{x},t)\Big{]}$
$+\sum_{l\in\mathcal{L}^{i}_{1_{u}}}\lambda_{l}\delta_{\mathbf{q}_{l},\mathbf{q}}\Big{[}\psi_{1_{u}}\big{(}\mathbf{q}^{+},\mathbf{x}^{+}\big{)}-\psi_{1_{u}}(\mathbf{q},\mathbf{x},t)\Big{]}$
$=$ $\sum_{l\in\mathcal{L}^{o}_{1_{u}}}\lambda_{l}\delta_{1_{u},\mathbf{q}}\Big{[}\mathbf{x}\mathbf{A}_{l}\delta_{1_{u},\mathbf{q}^{+}}-\mathbf{x}\delta_{1_{u},\mathbf{q}}\Big{]}$
$+\sum_{l\in\mathcal{L}^{i}_{1_{u}}}\lambda_{l}\delta_{\mathbf{q}_{l},\mathbf{q}}\Big{[}\mathbf{x}\mathbf{A}_{l}\delta_{1_{u},\mathbf{q}^{+}}-\mathbf{x}\delta_{1_{u},\mathbf{q}}\Big{]}$
$=$ $-\mathbf{x}\delta_{1_{u},\mathbf{q}}\sum_{l\in\mathcal{L}^{o}_{1_{u}}}\lambda_{l}+\sum_{l\in\mathcal{L}^{i}_{1_{u}}}\lambda_{l}\mathbf{x}\mathbf{A}_{l}\delta_{\mathbf{q}_{l},\mathbf{q}},$
where $\mathbf{q}_{l}$ is defined as the states directly connected to $1_{u}$, and $\mathbf{q}$ and $\mathbf{q}^{+}$ are states right before and after transitions at time $t$. Note that the second equality follows from (7), and the last equality follows because the self-transition has no contribution in both transition sets, which is the same for all states in our cases. Combined with (36), we have
$(L\psi)_{1_{u}}=k_{2}\mathbf{x}\delta_{1_{u},\mathbf{q}}-\mathbf{x}\delta_{1_{u},\mathbf{q}}\sum_{l\in\mathcal{L}^{o}_{1_{u}}}\lambda_{l}+\sum_{l\in\mathcal{L}^{i}_{1_{u}}}\lambda_{l}\mathbf{x}\mathbf{A}_{l}\delta_{\mathbf{q}_{l},\mathbf{q}}.$ (38)
Therefore, the expected value of the extended generator is
$\mathrm{E}\big{[}(L\psi)_{1_{u}}\big{]}=\Big{[}k_{2}-\sum_{l\in\mathcal{L}^{o}_{1_{u}}}\lambda_{l}\Big{]}\mathbf{v}_{1_{u}}+\sum_{l\in\mathcal{L}^{i}_{1_{u}}}\lambda_{l}\mathbf{v}_{\mathbf{q}_{l}}\mathbf{A}_{l}.$ (39)
We can then substitute (39) into Dynkin formula (10), which results in
$\frac{\mathrm{d}\mathrm{E}[\psi_{1_{u}}(\mathbf{q},\mathbf{x},t)]}{\mathrm{d}t}=\Big{[}k_{2}-\sum_{l\in\mathcal{L}^{o}_{1_{u}}}\lambda_{l}\Big{]}\mathbf{v}_{1_{u}}+\sum_{l\in\mathcal{L}^{i}_{1_{u}}}\lambda_{l}\mathbf{v}_{\mathbf{q}_{l}}\mathbf{A}_{l}.$ (40)
Assuming that the stability condition is satisfied, the derivative at the left-hand side shall equal to zero as $t\to\infty$. Hence, we obtain
$\mathbf{v}_{1_{u}}\Big{[}\sum_{l\in\mathcal{L}^{o}_{1_{u}}}\lambda_{l}-k_{2}\Big{]}=\sum_{l\in\mathcal{L}^{i}_{1_{u}}}\lambda_{l}\mathbf{v}_{\mathbf{q}_{l}}\mathbf{A}_{l},$ (41)
and the proof is completed.
Appendix B Derivations of Corollary 1: The Linear Hierarchy over Noisy Channel
Since we are only interested in the average AoII at the monitor, it is sufficient to track the expected value of the second entry of $\mathbf{v}_{\overline{\mathbf{q}}}$, namely, $v_{\overline{\mathbf{q}}}^{2}$ at each state $\overline{\mathbf{q}}$. Therefore, it turns out that (13) becomes a set of linear equations,
$\lambda v^{2}_{0_{e}}$ $=m_{1}\pi_{0_{e}}+p_{e}\mu v^{2}_{1_{m}}+p_{e}\mu v^{2}_{1_{u}},$ (42)
$(p\lambda+\mu)v^{2}_{1_{m}}$ $=\pi_{1_{m}}+p\lambda v^{2}_{0_{c}},$
$\mu v^{2}_{1_{u}}$ $=m_{2}\pi_{1_{u}}+\lambda v^{2}_{0_{e}}+p\lambda v^{2}_{1_{m}},$
where $v^{2}_{0_{c}}$ equals to zero since the extended generator $(L\phi)_{0_{c}}$ is always zero. We reformulate the above equations into the matrix form $\mathbf{B}\mathbf{v}^{\top}=\mathbf{b}^{\top}_{1}$, i.e.,
$\begin{bmatrix}1&-p_{e}\rho^{-1}&-p_{e}\rho^{-1}\\ 0&1&0\\ -\rho&-p\rho&1\end{bmatrix}\begin{bmatrix}v^{2}_{0_{e}}\\ v^{2}_{1_{m}}\\ v^{2}_{1_{u}}\end{bmatrix}=\begin{bmatrix}\frac{m_{1}\pi_{0_{e}}}{\lambda}\\ \frac{\pi_{1_{m}}}{p\lambda+\mu}\\ \frac{m_{2}\pi_{1_{u}}}{\mu}\end{bmatrix},$ (43)
where $\mathbf{b}^{\top}_{1}$ is the constant column vector at the right-hand side of (43). Hence, the solutions of the above linear equations are easy to obtain, that is,
$v^{2}_{0_{e}}$ $=\frac{m_{1}pp_{e}}{a\lambda}+p_{e}\rho^{-1}(v^{2}_{1_{m}}+v^{2}_{1_{u}}),$ (44)
$v^{2}_{1_{m}}$ $=\frac{pp_{c}}{a\lambda(p+\rho^{-1})^{2}},$
$v^{2}_{1_{u}}$ $=\frac{1}{p_{c}\mu}\Big{[}\frac{m_{2}(p^{2}\rho+pp_{e})}{a(p+\rho^{-1})}+\frac{m_{1}pp_{e}}{a}+\frac{pp_{c}(p_{e}\rho^{-1}+p)}{a(p+\rho^{-1})^{2}}\Big{]},$
Substituting the above results into (15), we have (17).
Appendix C Derivations of Corollary 2: The Quadratic Hierarchy over Noisy Channel
By leveraging (13) and (14) at each state, we obtain the set of linear equations of the quadratic case, i.e.,
$(\lambda-k_{1})v^{2}_{0_{e}}$ $=p_{e}\mu v^{2}_{1_{m}}+p_{e}\mu v^{2}_{1_{u}},$ (45)
$(p\lambda+\mu)v^{2}_{1_{m}}$ $=\pi_{1_{m}}+p\lambda v^{2}_{0_{c}},$
$(\mu-k_{2})v^{2}_{1_{u}}$ $=\lambda v^{2}_{0_{e}}+p\lambda v^{2}_{1_{m}},$
where $v^{2}_{0_{c}}$ equals to zero. We again reformulate the above equations into the matrix form $\mathbf{C}\mathbf{v}^{\top}=\mathbf{b}^{\top}_{2}$, i.e.,
$\begin{bmatrix}1&\frac{p_{e}\mu}{k_{1}-\lambda}&\frac{p_{e}\mu}{k_{1}-\lambda}\\ 0&1&0\\ \frac{\lambda}{k_{2}-\mu}&\frac{p\lambda}{k_{2}-\mu}&1\end{bmatrix}\begin{bmatrix}v^{2}_{0_{e}}\\ v^{2}_{1_{m}}\\ v^{2}_{1_{u}}\end{bmatrix}=\begin{bmatrix}0\\ \frac{\pi_{1_{m}}}{p\lambda+\mu}\\ 0\end{bmatrix}.$ (46)
Assuming that the stability conditions are satisfied where the inverse $\mathbf{C}^{-1}$ exists, the average AoII $x_{b}$ can be calculated by the summation of solutions of (46), that is,
$x_{b}=\mathbf{1}_{3}\mathbf{C}^{-1}\mathbf{b}^{\top}_{2},$ (47)
where $\mathbf{1}_{3}=[1,1,1]$ and $\mathbf{b}^{\top}_{2}$ is the constant column vector at the right-hand side of (46).
Appendix D Stability Conditions of the Quadratic Hierarchy over Noisy Channel
D-1 Proof of Necessity
We first prove the necessity of these two conditions. Given that the joint state $(\mathbf{q},\mathbf{x})$ is ergodic, the quadratic hierarchy AoII $x_{2}(t)$ converges to a limit $x_{2}$ that can be obtained by (47). Therefore, we are able to establish the necessary conditions from two aspects: 1) the invertibility of matrix $\mathbf{C}$; 2) the positivity of $v^{2}_{\overline{\mathbf{q}}}$ at each state $\overline{\mathbf{q}}$.
Following 1) and 2), we need to calculate the related quantities from (46). The determinant of $\mathbf{C}$ is obtained by the definition
$\det(\mathbf{C})=\sum_{j=1}^{3}c_{2j}M_{2j}=1-\frac{p_{e}\lambda\mu}{(k_{1}-\lambda)(k_{2}-\mu)}\neq 0,$ (48)
where $\{c_{2j}:j=1,2,3\}$ is the elements in the second row of $\mathbf{C}$ and $M_{2j}$ is the cofator of the element $c_{2j}$. Furthermore, the expected value $v^{2}_{\overline{\mathbf{q}}}$ of AoII can be calculated from (46),
$\begin{bmatrix}v^{2}_{0_{e}}\\ v^{2}_{1_{m}}\\ v^{2}_{1_{u}}\end{bmatrix}=\mathbf{C}^{-1}\mathbf{b}^{\top}_{2}=\frac{\pi_{1_{m}}}{p\lambda+\mu}\begin{bmatrix}(\mathbf{C}^{-1})_{12}\\ (\mathbf{C}^{-1})_{22}\\ (\mathbf{C}^{-1})_{32}\end{bmatrix},$ (49)
where the second column of $\mathbf{C}^{-1}$ can be obtained using the general formula for inverse matrix in linear algebra, i.e.,
$(\mathbf{C}^{-1})_{12}$ $=\frac{M_{21}}{\det(\mathbf{C})}=\frac{\frac{p_{e}\mu}{\lambda-k_{1}}+\frac{pp_{e}\lambda\mu}{(k_{1}-\lambda)(k_{2}-\mu)}}{\det(\mathbf{C})},$ (50)
$(\mathbf{C}^{-1})_{22}$ $=\frac{M_{22}}{\det(\mathbf{C})}=1,$
$(\mathbf{C}^{-1})_{32}$ $=\frac{M_{23}}{\det(\mathbf{C})}=\frac{\frac{p\lambda}{\mu-k_{2}}+\frac{p_{e}\lambda\mu}{(k_{1}-\lambda)(k_{2}-\mu)}}{\det(\mathbf{C})}.$
As a result, we only need to discuss the relationships between two parameter pairs, namely, $(\lambda,k_{1})$ and $(\mu,k_{2})$, that ensure the invertibility as well as the positivity of $(\mathbf{C}^{-1})_{12}$ and $(\mathbf{C}^{-1})_{32}$, which are divided into two cases.
- Case I: $\det(\mathbf{C})>0$ and $M_{21}>0$ and $M_{23}>0$. In this case, we shall discuss the following four situations.
I1). If $0<k_{1}<\lambda$, $0<k_{2}<\mu$, then we have
$(k_{1}-\lambda)(k_{2}-\mu)>p_{e}\lambda\mu,$ (51a)
$\mu-k_{2}+p\lambda>0,$ (51b)
$p(\lambda-k_{1})+p_{e}\mu>0,$ (51c)
where (51a), (51b) and (51c) follow from $\det(\mathbf{C})>0$, $M_{12}>0$ and $M_{23}>0$, respectively. The latter two equations can be further simplified as
$k_{2}<\mu+p\lambda,$ (52a)
$k_{1}<\lambda+\frac{p_{e}\mu}{p},$ (52b)
both of which satisfy the assumptions in I1).
I2). If $k_{1}>\lambda$, $k_{2}>\mu$, then we again have (51a), (51b) and (51c). However, the requirement in (51a) cannot be met since
$\max_{k_{1},k_{2}}(k_{1}-\lambda)(k_{2}-\mu)<\big{[}(\lambda+\frac{p_{e}\mu}{p})-\lambda\big{]}\big{[}(p\lambda+\mu)-\mu\big{]}=p_{e}\lambda\mu,$ (53)
where the first inequality follows from (52a) and (52b) that violates the condition $\det(\mathbf{C})>0$.
I3). If $0<k_{1}<\lambda$, $k_{2}>\mu$, then the condition $M_{23}>0$ cannot be satisfied since it yields
$p(\lambda-k_{1})+p_{e}\mu<0,$ (54)
which requires $k_{1}>\lambda+\frac{p_{e}\mu}{p}$.
I4). If $k_{1}>\lambda$, $0<k_{2}<\mu$, then the condition $M_{21}>0$ cannot be satisfied since it yields
$\mu-k_{2}+p\lambda<0,$ (55)
which requires $k_{2}>\mu+p\lambda$.
Next, we consider the other case.
- Case II: $\det(\mathbf{C})<0$ and $M_{21}<0$ and $M_{23}<0$. We also need to discuss the same four situations as case I.
II1). If $0<k_{1}<\lambda$, $0<k_{2}<\mu$, then we have
$(k_{1}-\lambda)(k_{2}-\mu)<p_{e}\lambda\mu,$ (56a)
$\mu-k_{2}+p\lambda<0,$ (56b)
$p(\lambda-k_{1})+p_{e}\mu<0,$ (56c)
where (56a), (56b) and (56c) follow from $\det(\mathbf{C})<0$, $M_{12}<0$ and $M_{23}<0$, respectively. The latter two equations can be further simplified as
$k_{2}>\mu+p\lambda,$ (57a)
$k_{1}>\lambda+\frac{p_{e}\mu}{p},$ (57b)
both of which violate the assumptions in II1).
II2). If $k_{1}>\lambda$, $k_{2}>\mu$, then we again have (56a), (56b) and (56c). However, the requirement in (56a) cannot be met since
$\min_{k_{1},k_{2}}(k_{1}-\lambda)(k_{2}-\mu)>\big{[}(\lambda+\frac{p_{e}\mu}{p})-\lambda\big{]}\big{[}(p\lambda+\mu)-\mu\big{]}=p_{e}\lambda\mu,$ (58)
where the first inequality follows from (57a) and (57b) that violates the condition $\det(\mathbf{C})<0$.
II3). If $0<k_{1}<\lambda$, $k_{2}>\mu$, then the condition $\det(\mathbf{C})<0$ cannot be satisfied since
$\det(\mathbf{C})=1-\frac{p_{e}\lambda\mu}{(k_{1}-\lambda)(k_{2}-\mu)}>0,$ (59)
under our assumptions in II3) due to the negativity of $(k_{1}-\lambda)(k_{2}-\mu)$.
II4). If $k_{1}>\lambda$, $0<k_{2}<\mu$, then the condition $\det(\mathbf{C})<0$ cannot be satisfied for the same reason as II3).
To summarize, we have proved the necessity of the stability conditions consisting of $0<k_{1}<\lambda$, $0<k_{1}<k_{2}<\mu$ and (51a) as given in Theorem 2.
D-2 Proof of Sufficiency
In order to prove the sufficiency of these two conditions, we need to give the ergodicity of the joint state $(\mathbf{q},\mathbf{x})$ also the boundness of the average AoII $x_{b}$. Let $(\mathbf{q}_{0},\mathbf{x}_{0})$ denotes the initial condition at time $t_{0}$ and $\{t_{i}:i\geqslant 0\}$ be the sequence of transition time. On each interval $[t_{i},t_{i+1})$, the AoII at the monitor evolves as
$x_{2}(s)=x_{2}(t_{i})+\int_{t_{i}}^{s}f_{2}^{\prime}(\mathbf{q}(r),\mathbf{x}(r),r)\mathrm{d}r,\ \forall s\in[t_{i},t_{i+1}),$ (60)
according to (3). Taking norms at both sides we have
$||x_{2}(s)||\leqslant||x_{2}(t_{i})||+\int_{t_{i}}^{s}\max\{k_{2}||x_{2}(r)||,c\}\mathrm{d}r,$ (61)
where $c=1$ is the given slope constant at linear growth in our case [^8] It is easy to verify that $c$ can be any positive constant., and we assume that $k_{2}>k_{1}$ without loss of generality. Since $k_{2}||x_{2}(r)||$ is continuous on each interval $[t_{i},t_{i+1})$, one of the following cases must occur:
I1). $k_{2}||x_{2}(r)||<1$. Then we have
$||x_{2}(s)||\leqslant||x_{2}(t_{i})||+t_{i+1}-t_{i}.\forall s\in[t_{i},t_{i+1}),$ (62)
I2). $k_{2}||x_{2}(r)||>1$. Then we have
$||x_{2}(s)||$ $\leqslant||x_{2}(t_{i})||+\int_{t_{i}}^{s}k_{2}||x_{2}(r)||\mathrm{d}r$ (63)
$\leqslant\exp\Big{[}\int_{t_{i}}^{s}k_{2}\mathrm{d}r\Big{]}||x_{2}(t_{i})||,\forall s\in[t_{i},t_{i+1}),$
where the second inequality follows from the Bellman-Gronwall Lemma.
I3). $k_{2}||x_{2}(r_{0})||=1$ for some $r_{0}\in[t_{i},t_{i+1})$. Then we can leverage the Bellman-Gronwall Lemma on $r\in[r_{0},t_{i+1})$ to obtain
$||x_{2}(s)||$ $\leqslant\exp\Big{[}\int_{r_{0}}^{s}k_{2}\mathrm{d}r\Big{]}||x_{2}(r_{0})||,\ \forall s\in[r_{0},t_{i+1}).$ (64)
Combining I2) and I3), for $\forall s\in[t_{i},t_{i+1})$ we have
$||x_{2}(s)||\leqslant\exp\Big{[}\int_{t_{i}}^{s}k_{2}\mathrm{d}r\Big{]}\max\{||x_{2}(t_{i})||,\frac{1}{k_{2}}\}.$ (65)
Considering the transition map $\phi_{\mathit{l}}$ defined in (4) at time $t_{i+1}$, we have
$||x_{2}(t_{i+1})||$ $=||\phi^{x}_{l}(\mathbf{\mathbf{q},\mathbf{x},t})||$ (66)
$\leqslant\exp\Big{[}\int_{t_{i}}^{t_{i+1}}k_{2}\mathrm{d}r\Big{]}\max\{||x_{2}(t_{i})||,\frac{1}{k_{2}},0\},$
where $\phi^{x}_{l}$ is the mapping of the continuous state. The inequality follows from Table I since the AoII at the monitor either keeps growing or drops to zero at the transition time $t_{i+1}$.
In the worst case, the second entry of initial condition $\mathbf{x}_{0}=(x_{1}(t_{0}),x_{2}(t_{0}))$ is larger than the constant $\frac{1}{k_{2}}$ and the first term in the maximum function of (64) always dominates as the AoII grows from $i=0$ to $i=k-1$, which yields
$||x_{2}(t_{k})||\leqslant\exp\Big{[}\int_{t_{0}}^{t_{k}}k_{2}\mathrm{d}r\Big{]}||x_{2}(t_{0})||.$ (67)
Denoting $t_{k}$ as the last transition time before some arbitrary time $t>t_{k}$, we obtain
$||x_{2}(t)||$ $\leqslant\exp\Big{[}\int_{t_{k}}^{t}k_{2}\mathrm{d}r\Big{]}\max\{||x_{2}(t_{k})||,\frac{1}{k_{2}}\}$ (68)
$\leqslant\exp\Big{[}\int_{t_{0}}^{t}k_{2}\mathrm{d}r\Big{]}||x_{2}(t_{0})||,\forall t\in[t_{k},t_{k+1}),$
from (64) and (67) for $i=k$ and $s=t$. Note that we have ergodicity of discrete state $\mathbf{q}\in\mathbb{Q}$, which implies the state $(\mathbf{q},x_{2}(t))$ would eventually return to $(0_{c},0)$ at some $t_{i+1}$ in a finite time even as $t\to\infty$. Therefore, $||x_{2}(t)||$ is always bounded in the quadratic hierarchy case, which in turn gives the ergodicity of the joint state $(\mathbf{q},\mathbf{x})$ as well as the boundness of the average AoII $x_{b}$.
Appendix E Derivations of Corollary 3: The Quadratic Hierarchy over Collision Channel
By leveraging (13) and (14) at each state $\mathbf{q}\in\mathbb{Q}_{c}$, we obtain the set of linear equations, i.e.,
$(p\lambda+\lambda_{c}+\mu)v^{2}_{1_{m}}$ $=\pi^{\prime}_{1_{m}}+\mu_{c}v^{2}_{2_{m}},$ (69)
$(p\lambda+\mu+\mu_{c})v^{2}_{2_{m}}$ $=\pi^{\prime}_{2_{m}}+\lambda_{c}v^{2}_{1_{m}},$
$(\lambda+\lambda_{c}-k_{1})v^{2}_{0_{e}}$ $=\mu_{c}v^{2}_{1_{e}},$
$(\lambda+\mu_{c}-k_{1})v^{2}_{1_{e}}$ $=\mu v^{2}_{2_{m}}+\lambda_{c}v^{2}_{0_{e}}+\mu v^{2}_{2_{u}},$
$(\lambda_{c}+\mu-k_{1})v^{2}_{1_{u}}$ $=p\lambda v^{2}_{1_{m}}+\lambda v^{2}_{0_{e}}+\mu_{c}v^{2}_{2_{u}},$
$(\mu+\mu_{c}-k_{2})v^{2}_{2_{u}}$ $=p\lambda v^{2}_{2_{m}}+\lambda v^{2}_{1_{e}}+\lambda_{c}v^{2}_{1_{u}},$
where $v^{2}_{0_{c}}=v^{2}_{1_{c}}=0$ is omitted above since $0_{c}$ and $1_{c}$ are irrelevant states. Note that the expected AoII at states $1_{m}$ and $2_{m}$ grows like the traditional AoI, which are only related to each other. These two equations can be solved first
$v^{2}_{1_{m}}=\frac{(p\lambda+\mu+\mu_{c})\pi_{1_{m}}^{\prime}+\mu_{c}\pi_{2_{m}}^{\prime}}{(p\lambda+\lambda_{c}+\mu)(p\lambda+\mu+\mu_{c})-\lambda_{c}\mu_{c}}=c_{1},$ (70)
$v^{2}_{2_{m}}=\frac{(p\lambda+\lambda_{c}+\mu)\pi_{2_{m}}^{\prime}+\lambda_{c}\pi_{1_{m}}^{\prime}}{(p\lambda+\lambda_{c}+\mu)(p\lambda+\mu+\mu_{c})-\lambda_{c}\mu_{c}}=c_{2},$
and denote as constant $c_{1}$ and $c_{2}$, respectively. Therefore, (69) can be reduced to a set of $4\times 4$ linear equations with the matrix form $\mathbf{D}\mathbf{v}^{\top}=\mathbf{b}^{\top}_{3}$, i.e.,
$\begin{bmatrix}1&\frac{\mu_{c}}{k_{1}-\lambda-\lambda_{c}}&0&0\\ \frac{\lambda_{c}}{k_{1}-\lambda-\mu_{c}}&1&0&\frac{\mu}{k_{1}-\lambda-\mu_{c}}\\ \frac{\lambda}{k_{1}-\lambda_{c}-\mu}&0&1&\frac{\mu_{c}}{k_{1}-\lambda_{c}-\mu}\\ 0&\frac{\lambda}{k_{2}-\mu-\mu_{c}}&\frac{\lambda_{c}}{k_{2}-\mu-\mu_{c}}&1\end{bmatrix}\begin{bmatrix}v^{2}_{0_{e}}\\ v^{2}_{1_{e}}\\ v^{2}_{1_{u}}\\ v^{2}_{2_{u}}\end{bmatrix}$ (71)
$=\begin{bmatrix}0\\ \frac{\mu c_{2}}{\lambda+\mu_{c}-k_{1}}\\ \frac{p\lambda c_{1}}{\lambda_{c}+\mu-k_{1}}\\ \frac{p\lambda c_{2}}{\mu+\mu_{c}-k_{2}}\end{bmatrix}.$
Assuming that the stability conditions are satisfied where the inverse $\mathbf{D}^{-1}$ exists, the average AoII $x_{c}$ can be calculated by the summation of solutions of (70) and (71), that is,
$x_{c}=c_{1}+c_{2}+\mathbf{1}_{4}\mathbf{D}^{-1}\mathbf{b}^{\top}_{3},$ (72)
where $\mathbf{1}_{4}=[1,1,1,1]$ and $\mathbf{b}^{\top}_{3}$ is the constant column vector at the right-hand side of (71).
Appendix F Sufficient Stability Conditions of the Quadratic Hierarchy over Collision Channel
The step to prove the ergodicity is the same as the proof 2) in Appendix D and is omitted here. The determinant of $\mathbf{D}$ is obtained in (73) using the definition.
The expected values of AoII at states $1_{m},2_{m}$ are given in (70) as $c_{1}$ and $c_{2}$, both of which are larger than zero. Then we need to calculate the remaining expected values of AoII at states $0_{e},1_{e},1_{u},2_{u}$, which are summarized in (74).
$\det(\mathbf{D})=$ $\frac{\Pi}{(\lambda+\mu_{c}-k_{1})(\lambda+\lambda_{c}-k_{1})(\lambda_{c}+\mu-k_{1})(\mu+\mu_{c}-k_{2})},$ (73)
[…]
$+\big{[}\mu(m-k_{2})+(\mu+\mu_{c})(\lambda_{c}+\mu_{c}-k_{2})+2\lambda(\mu_{c}-k_{2})-2k_{2}\lambda_{c}\big{]}k^{2}_{1}$
$-\big{[}(\lambda_{c}+\mu+\mu_{c}-k_{2})(\lambda\mu+\lambda\mu_{c}+\lambda_{c}\mu+\mu\mu_{c})+\lambda(\lambda+\mu)(\mu_{c}-k_{2})-3k_{2}\lambda\lambda_{c}\big{]}k_{1}\biggr{\}}.$
$v^{2}_{0_{e}}$ $=\frac{1}{\Pi}\biggl{\{}c_{2}\mu\mu_{c}A+c_{1}p\lambda\lambda_{c}\mu\mu_{c}\biggr{\}},$ (74)
$v^{2}_{1_{e}}$ $=\frac{1}{\Pi}\biggl{\{}c_{2}\mu(\lambda+\lambda_{c}-k_{1})A+c_{1}p\lambda\lambda_{c}\mu(\lambda+\lambda_{c}-k_{1})\biggr{\}},$
$v^{2}_{1_{u}}$ $=\frac{1}{\Pi}\biggl{\{}c_{2}\lambda\mu\mu_{c}(m-k_{1}-k_{2})+(c_{2}p\lambda\mu_{c}+\frac{c_{1}p\lambda\lambda_{c}\mu_{c}}{\lambda_{c}+\mu-k_{1}})B+\frac{c_{1}p\lambda\Pi}{\lambda_{c}+\mu-k_{1}}\biggl{\}},$
$v^{2}_{2_{u}}$ $=\frac{1}{\Pi}\biggl{\{}c_{2}\lambda\mu\big{[}\lambda_{c}(m-2k_{1})+(\lambda-k_{1})(\mu-k_{1})\big{]}+p\lambda(\lambda-k_{1})(\lambda+\lambda_{c}+\mu_{c}-k_{1})\big{[}c_{1}\lambda_{c}+c_{2}(\lambda_{c}+\mu-k_{1})\big{]}\biggr{\}},$
$A$ $=(k_{2}-p\lambda-\mu-\mu_{c})k_{1}-(\lambda_{c}+\mu)k_{2}+(\lambda_{c}+\mu+\mu_{c})\mu+(\lambda_{c}+\mu)p\lambda,$
$B$ $=k_{1}^{2}-(2\lambda+\lambda_{c}+\mu_{c})k_{1}+m\lambda.$
$\Pi^{\prime}$ $=(\lambda+\mu-k_{1})C,$ (75)
$C$ $=2\mu^{2}(\lambda-k_{1})+2k^{2}_{1}\mu-(4\lambda-k_{2})k_{1}\mu-(2\lambda-k_{1})(\lambda-k_{1})k_{2}-k_{2}\lambda\mu$
With the above quantities prepared, we are now ready to verify our sufficient conditions proposed in Theorem 3 by checking the positivity of (74) and the invertibility of $\mathbf{D}$. Given that $0<k_{1}<\lambda<\mu$ and $0<k_{2}<\mu$, we only need to verify the positivity of $A$ and $B$ defined in (74) since the rest parts are easily checked to be positive. Consequently, we have the following discussions.
I). Since $A$ is a monotonically decreasing linear function of $k_{1}$, we have
$\min_{k_{1}}A$ $>(k_{2}-p\lambda-\mu-\mu_{c})\lambda-(\lambda_{c}+\mu)k_{2}$ (76)
$+(\lambda_{c}+\mu+\mu_{c})\mu+(\lambda_{c}+\mu)p\lambda$
$=(\lambda-\lambda_{c}-\mu)k_{2}+(\lambda_{c}+\mu+\mu_{c})\mu$
$+(\lambda_{c}+\mu)p\lambda-(p\lambda+\mu+\mu_{c})\lambda$
$=A^{\prime}$
where $A^{\prime}$ is again a monotonically decreasing linear function of $k_{2}$. Given $k_{2}<\mu$, we have
$A^{\prime}$ $>(\lambda-\lambda_{c}-\mu)\mu+(\lambda_{c}+\mu+\mu_{c})\mu$ (77)
$+(\lambda_{c}+\mu)p\lambda-(p\lambda+\mu+\mu_{c})\lambda$
$=\mu\mu_{c}+(\lambda_{c}+\mu)p\lambda-(p\lambda+\mu_{c})\lambda$
$=(\mu-\lambda)(p\lambda+\mu_{c})+p\lambda\lambda_{c}$
$>0.$
Hence, we can obtain
$A>\min_{k_{1}}A>A^{\prime}>0,$ (78)
which guarantees the positivity of constant $A$.
II). Constant $B$ is a parabola of $k_{1}$, whose direction of opening is upward. The focus of this parabola is at point $(\lambda+\frac{\lambda_{c}+\mu_{c}}{2},m\lambda-(\frac{\lambda_{c}+\mu_{c}}{2})^{2})$. However, since the curve of $B$ is limited by the value of $k_{1}$, we can obtain the minimal of $B$ in the interval $k_{1}\in(0,\lambda)$, i.e.,
$\min_{k_{1}}B>\lambda^{2}-(2\lambda+\lambda_{c}+\mu_{c})\lambda+m\lambda=\lambda\mu>0,$ (79)
where the equality follows from $m=\lambda+\lambda_{c}+\mu+\mu_{c}$. Hence, we obtain the positivity of constant $B$.
Since we have the positivity of (74), combined with $\Pi$ in (73) being positive, we obtain the sufficient stability conditions of the general case in Theorem 3.
Now, we can have a finer result of condition i) in Theorem 3 for the symmetric case where $\lambda=\lambda_{c}$ and $\mu=\mu_{c}$. To that end, we first simplify $\Pi$ as $\Pi^{\prime}$ in (75). Since we have
$2\lambda-k_{1}>$ $0,$ (80)
$2\mu-k_{2}>$ $0,$
$\lambda+\mu-k_{1}>$ $0,$
using condition ii), we only need to guarantee the constant $C$ to be positive in order to satisfy condition i) under the symmetric case. As a result, we have
$C=$ $Dk_{2}+2k^{2}_{1}\mu+2\lambda\mu^{2}-4k_{1}\lambda\mu-2k_{1}\mu^{2},$ (81)
$D=$ $-k_{1}^{2}+(3\lambda+\mu)k_{1}-\lambda\mu-2\lambda^{2}$
Again, we have a parabola $D$ of $k_{1}$ whose direction of opening is downward. Since the curve of this parabola is limited by $k_{1}\in(0,\lambda)$, the maximal of $D$ is less than the value when $k_{1}=\lambda$, i.e.,
$\max_{k_{1}\in(0,\lambda)}D<D_{k_{1}=\lambda}=0.$ (82)
To ensure $\det(\mathbf{D})>0$, we need $C>0$, which results in
$k_{2}<\frac{2k_{1}\mu^{2}+4k_{1}\lambda\mu-2k_{1}^{2}\mu-2\lambda\mu^{2}}{(3\lambda+\mu)k_{1}-k_{1}^{2}-\lambda\mu-2\lambda^{2}}.$ (83)
Combined with the above results, we have completed the proof of Theorem 3.